OpenStudy (anonymous):

\left(\begin{matrix}t \\ -19\end{matrix}\right) > -5

OpenStudy (anonymous):

\[\left(\begin{matrix}t \\ -19\end{matrix}\right) > -5\]

OpenStudy (anonymous):

how do I simplify this ? and work out the answer ?

OpenStudy (anonymous):

this doesn't make sense. How can a vector be bigger than a number?

OpenStudy (anonymous):

maybe you mean this: \[\left(\begin{matrix}t \\ -19\end{matrix}\right)\left(\begin{matrix}x \\ y\end{matrix}\right)>-5\]

OpenStudy (anonymous):

Can you explain what you are trying to do?

OpenStudy (anonymous):

i mean should be reverse and X^T, just don't know how to write that....

OpenStudy (anonymous):

the book just simply says solved \[\left(\begin{matrix}t \\ -19\end{matrix}\right) > -5\] o.0

OpenStudy (anonymous):

draw the problem.

OpenStudy (anonymous):